# AED-PADA:Improving Generalizability of Adversarial Example Detection via Principal Adversarial Domain Adaptation

### v1

@@ 4.1 Experimental Setups
To evaluate the generalization capabilities of the detection methods, it is important to consider a diverse set of attack methods, including both the earlier and recent techniques. Here, 10 earlier attack methods are utilized to generate adversarial examples for training, including FGSM [?], BIM [?], C&W [?], DeepFool [?], PGD [?], MIFGSM [?], DIM[?], ILA [?], YA-ILA [?] and SINIFGSM [?]. 7 SOTA attack methods are employed to generate adversarial examples for testing, including APGD [?], ILA-DA [?], Jitter [?], SSA [?], TIFGSM [?], VMIFGSM [?] and VNIFGSM [?].
Implementation details.
In this paper, adversarial examples are generated to carry out the untargeted white-box attack with the $l_{\infty}$ norm constraint. The maximum magnitude of adversarial perturbation is set to 2. Step size and number of iterations for adversarial attacks are set to 1/255 and 10, respectively. During PADA stage, we utilize MFSAN [?] as the basic MDA method, and employ the same training strategy as it. All the adversarial example detection methods are trained consistently for 100 epochs. To evaluate the performance of our proposed method, the widely used Accuracy is employed as the metric for the adversarial example detection task. The experiments are conducted on an NVIDIA GeForce RTX 3080Ti GPU.
More details of the experimental setups can be found in the supplementary materials.
4.2 Performance Evaluations
Here, we compare our method with 5 SOTA detection methods, including LID [?], MD [?], Steg [?], SID [?] and SA [?].
Performances of defending unseen attacks.
Table 1 presents the cross-attack detection results on 7 unseen state-of-the-art testing adversarial attacks. These results provide compelling evidence that our approach obviously outperforms the SOTA methods in terms of the generalization ability. Note that the performance of our approach (91.219%) surpasses that of MD-PADs (52.266%) by a significant margin, i.e., 38.953%, under the settings of SVHN and VGG-16. Besides, PADs may not be suitable to be directly applied to the existing methods. As can be observed, if the SOTA detection methods directly adopt PADs for training (denoted as X-PADs), their performances may not always increase. This observation further verifies the effectiveness of our PADA in our AED-PADA framework.
@@ 4.2 Performance Evaluations
APGD ILA-DA Jitter SSA TIFGSM VMIFGSM VNIFGSM
CIFAR-10 (ResNet-18) LID 90.044 94.121 78.240 59.839 90.863 91.056 86.496 84.380
@@
SA 84.907 90.731 87.365 81.323 89.275 90.013 91.967 87.940
SA-PADs 89.125 93.485 91.925 88.885 94.76 94.875 95.595 92.664
AED-PADA (Ours) 90.545 97.855 97.065 84.255 97.700 97.765 97.675 94.694
CIFAR-10 (VGG-16) LID 85.024 93.357 75.092 58.639 88.201 88.173 81.540 81.432
@@
AED-PADA (Ours) 84.370 94.365 93.260 65.765 93.135 93.410 93.545 88.264
SVHN (ResNet-18) LID 66.431 87.940 63.697 60.324 71.403 71.384 64.952 69.447
@@
AED-PADA (Ours) 73.500 99.220 98.995 68.990 99.295 99.310 99.220 91.219
@@
Table 1: Comparison of the generalization performances between the state-of-the-art adversarial detection methods and our AED-PADA. For the methods without the ‘-PADs’ mark, since they are trained on single attack, each of their results is the averaged detection accuracy with 10 training attacks being utilized one after another in training. The methods with the ‘-PADs’ mark are trained by the adversarial examples in PADs, which are consistent with our AED-PADA.
Generalization performances across different backbones and datasets.
@@
To thoroughly evaluate the generalization ability of adversarial detection methods, it is also vital to assess their performances in the scenarios where the training and testing adversarial examples are from different backbones and datasets. A detection method with better cross-attack, cross-backbone, and cross-dataset capabilities tends to be better suited for practical applications in complex environments. Table 2 displays the generalization performances of different detection methods across different backbones and datasets. In general, results from both scenarios demonstrate that our method exhibits the best generalization ability. Since the structures tend to vary significantly among different types of backbones, which will result in different feature dimensions between training and testing, both LID and MD are not applicable in cross-backbone experiment, i.e., they lack the capability to generalize across backbones, because they utilize the intermediate features of the backbones for detection.
@@
Backbone Dataset
@@
Table 2: Comparison of generalization performances across different backbones and datasets. The cross-backbone results are obtained on CIFAR-10, and the cross-dataset results are obtained when the backbone is ResNet-18.
@@
Same Cluster FGSM+SINIFGSM+PGD 90.444 -
@@
Table 3: Comparison of generalization performances when different source domains are employed in the training process. The dataset is CIFAR-10, and the backbone is ResNet-18. The result of AD clustering is {FGSM, PGD, DIM, MIFGSM, SINIFGSM}-{BIM, ILA, YA-ILA}-{C&W, DeepFool}. ‘Same Cluster’ and ‘Cross Cluster’ respectively represent that the ADs employed as the sources domains are selected from the same AD cluster and different AD clusters.
@@
CIFAR-10 (ResNet-18) 91.936 94.694 94.784 94.571 94.561 95.018
CIFAR-10 (VGG-16) 87.946 88.264 87.964 87.991 88.104 88.498
SVHN (ResNet-18) 91.497 94.344 91.257 91.726 94.999 94.801
SVHN (VGG-16) 89.183 89.269 91.219 90.056 90.026 90.713
Table 4: Results with different number of clusters $K$. The underlined and bolded values respectively represent the results of our AED-PADA, in which the number of clusters is automatically determined based on the CH score, and the best averaged accuracies.
@@
Table 5: Results of different feature extractions in PADA of our AED-PADA.
Dataset Backbone Averaged Accuracy (%)
MFSAN (Ours) M3DA-$\beta$
CIFAR-10 ResNet-18 94.694 94.976
VGG-16 88.864 87.819
SVHN ResNet-18 94.344 93.892
VGG-16 91.219 90.960
Table 6: Results of our framework with different MDA methods.
4.3 Analysis
Effectiveness of AD clustering and PADs Selection.
As shown in Table 3, we set up two groups of experiments, i.e., selecting ADs from Same Cluster and Cross Cluster, to verify the effectiveness of our AD clustering and PADs Selection, against the 7 unseen attacks. According to Table 3, for ‘Same Cluster’, the best and worst averaged results are 90.444% and 93.494%, respectively. For ‘Cross Cluster’, the best, worst, and averaged values are 94.694%, 92.429%, and 93.892%, respectively. Then, we can obtain three observations: (1) The averaged results of ‘Cross Cluster’ selection is higher than the best result of ‘Same Cluster’ selection, and the best result of ‘Cross Cluster’ selection is also superior. This verifies the effectiveness of our AD Clustering. Apparently, source domains from ‘Cross Cluster’ can certainly enhance the generalization ability of the detection methods. (2) Although ‘Cross Cluster’ selection in general achieves better results, it cannot guarantee that the randomly selected ADs (from ‘Cross Cluster’s) always give better performance than the results of ‘Same Cluster’ selected ADs. When the PADs with a very low CEFS score are employed as the source domains for detection, the generalization performance may be worse than that of ‘Same Cluster’ selection. This observation further verifies the effectiveness of our PADs Selection. (3) For the majority of results, a higher CEFS value of PADs induces a higher averaged accuracy of the detection method across different target domains. This observation suggests that the proposed CEFS is a proper guide for selecting PADs, i.e., a higher CEFS score indicates that the corresponding PADs can cover a larger proportion of the entire feature space, and thus gives a better performance.
The effect of different numbers of clusters.
Here, we apply AD Clustering to 10 adversarial attack methods in the training set and present the impact of different cluster numbers $K$ (the number of adversarial attacks in PADs). As depicted in Table 4, the detection performance increase when $K$ increases, until a certain value of $K$ is achieved. This trend is attributed to the utilization of an increased number of ADs as the source domains, which undeniably provides larger coverage of the entire feature space. Note that we can employ the CH score to automatically determine the number of clusters, which is included in AD Clustering of our AED-PADA, and these results are underlined in Table 4. As can be observed, our automatic CH score based method can yield superior performance on SVHN with VGG-16 being the backbone, and give comparable performance to the best result on other settings. Although the best result with manually selecting $K$ can exceeds our automatic method for 0.2%-0.3% in terms of accuracy, on average, it requires 2.3×2.3$\times$2.3 × more parameters and gives a 3×3$\times$3 × slower speed. This suggests that our method actually achieves a better balance between the training costs and generalization performances.
Effectiveness of the adversarial feature enhancement in PADA.

### v2

@@ 4.1.4. Baseline adversarial attack methods
To evaluate the generalization capabilities of the detection methods, it is important to consider a diverse set of attack methods, including both the earlier and recent techniques. Here, 10 earlier attack methods are utilized to generate adversarial examples for training, including FGSM (FGSM, ), BIM (BIM, ), C&W (CW, ), DeepFool (DeepFool, ), PGD (pgd, ), MI-FGSM (MIM, ), DIM(DIM, ), ILA (ILA, ), YA-ILA (ILA++, ) and SI-NI-FGSM (si-ni-fgsm, ). 7 SOTA attack methods are employed to generate adversarial examples for testing, including APGD (APGD, ), ILA-DA (ILA-DA, ), Jitter (jitter, ), SSA (SSA, ), TI-FGSM (TIM, ), VMI-FGSM (vmifgsm, ) and VNI-FGSM (vmifgsm, ). The adversarial attacks for training and testing are entirely distinct, so the adversarial detection results on the unseen testing attacks indicate the generalization performance of our proposed detection method.
4.1.5. Baseline adversarial detection methods
We compare the generalizability of our AED-PADA with five state-of-the-art adversarial detection methods, LID (lid, ), MD (md, ), Steg (steg, ), SID (SID, ) and SA (txt_advdetection, ). These detection methods differ from ours as they employ only a single attack method for training. To ensure a fair comparison with our AED-PADA, it is necessary to consider the scenario training by multiple attacks. Consequently, we consider the following two configurations for the SOTA detection methods: (1) For single attack training, we utilize all data from Train-acq-src to ensure sufficient data volume for effective training. (2) For multiple attacks training, the methods are trained by the adversarial examples in PADs, which are consistent with our AED-PADA. Each source domain contains 10,000 samples from Train-acq-src, evenly divided into 5,000 benign and 5,000 adversarial examples.
Table 3. Comparison of the generalization performances on CIFAR-10 between the state-of-the-art adversarial detection methods and our AED-PADA. Averaged Accuracy is the average result across seven unseen state-of-the-art testing adversarial attacks. The bolded and the underlined values represent the best and the second best results for each column, respectively.
@@ 4.1.5. Baseline adversarial detection methods
APGD ILA-DA Jitter SSA TI-FGSM VMI-FGSM VNI-FGSM
CIFAR-10 (ResNet-18) LID (lid, ) 90.044 94.121 78.240 59.839 90.863 91.056 86.496 84.380
@@
SA (txt_advdetection, ) 84.907 90.731 87.365 81.323 89.275 90.013 91.967 87.940
SA-PADs 89.125 93.485 91.925 88.885 94.760 94.875 95.595 92.664
AED-PADA 90.545 97.855 97.065 84.255 97.700 97.765 97.675 94.694
CIFAR-10 (VGG-16) LID (lid, ) 85.024 93.357 75.092 58.639 88.201 88.173 81.540 81.432
@@
AED-PADA 84.370 94.365 93.260 65.765 93.135 93.410 93.545 88.264
Table 4. Comparison of the generalization performances on SVHN between the state-of-the-art adversarial detection methods and our AED-PADA. Averaged Accuracy is the average result across seven unseen state-of-the-art testing adversarial attacks. The bolded and the underlined values represent the best and the second best results for each column, respectively.
APGD ILA-DA Jitter SSA TI-FGSM VMI-FGSM VNI-FGSM
SVHN (ResNet-18) LID (lid, ) 66.431 87.940 63.697 60.324 71.403 71.384 64.952 69.447
@@
AED-PADA 73.500 99.220 98.995 68.990 99.295 99.310 99.220 91.219
Table 5. Comparison of the generalization performances on ImageNet between the state-of-the-art adversarial detection methods and our AED-PADA. Averaged Accuracy is the average result across seven unseen state-of-the-art testing adversarial attacks. The bolded and the underlined values represent the best and the second best results for each column, respectively.
APGD ILA-DA Jitter SSA TI-FGSM VMI-FGSM VNI-FGSM
ImageNet (ResNet-18) LID (lid, ) 52.868 54.078 50.991 53.263 53.793 56.566 56.549 54.015
LID-PADs 57.340 63.910 55.195 58.150 58.890 60.900 61.245 59.376
MD (md, ) 53.760 52.031 51.728 51.603 53.533 59.438 58.536 54.376
MD-PADs 54.560 50.425 50.685 52.765 54.695 62.165 61.040 55.191
Steg (steg, ) 78.415 86.954 86.553 85.431 65.101 94.679 94.606 84.534
Steg-PADs 90.725 94.735 94.700 94.535 89.660 94.875 94.875 93.444
SID (SID, ) 51.885 56.475 51.745 51.425 55.285 55.335 54.995 53.878
SID-PADs 55.070 56.975 53.785 52.335 59.790 59.860 58.400 56.602
SA (txt_advdetection, ) 85.738 87.149 85.912 85.712 86.960 85.988 86.014 86.210
SA-PADs 66.955 69.230 67.665 66.930 67.385 68.490 68.550 67.886
AED-PADA 98.830 99.935 99.965 99.965 99.070 99.965 99.960 99.670
ImageNet (VGG-16) LID (lid, ) 53.144 53.672 52.029 51.694 52.056 57.081 56.173 53.693
LID-PADs 56.295 59.330 55.110 54.825 55.695 59.960 59.310 57.218
MD (md, ) 53.760 52.031 51.728 51.603 53.533 59.438 58.536 54.376
MD-PADs 53.470 50.705 51.960 50.970 52.430 57.385 56.585 53.358
Steg (steg, ) 78.931 90.915 85.852 84.279 66.300 94.695 94.969 85.134
Steg-PADs 90.585 94.185 94.115 94.165 90.375 94.530 94.535 93.213
SID (SID, ) 41.718 50.624 50.477 44.102 51.015 51.567 51.528 48.719
SID-PADs 43.238 46.165 53.794 50.002 49.125 53.257 46.382 48.852
SA (txt_advdetection, ) 87.747 93.025 89.939 88.176 86.256 92.133 92.255 89.933
SA-PADs 91.590 94.910 93.415 92.310 91.280 94.250 94.280 93.148
AED-PADA 98.960 99.930 99.935 99.890 98.225 99.935 99.930 99.544
4.1.6. Implementation details
In this paper, adversarial examples are generated by the untargeted white-box attacks with the $l_{\infty}$ norm constraint. The perturbation of the training adversarial examples is constrained to a challenging scenario, where the maximum magnitude of the adversarial perturbation is set to 2. The step size and number of iterations for adversarial attacks are set to 1/255 and 10, respectively. During PADA stage, we utilize MFSAN (aaai-mda, ) as the basic MUDA method and employ the same training strategy as it. All the adversarial example detection methods are trained consistently for 100 epochs. To evaluate the performance of our proposed method, the widely used Accuracy is employed as the metric for the adversarial example detection task. We utilize MFSAN as the basic MUDA method in the PADA stage of our framework and set the trade-off parameters $\lambda=\gamma=1.0$, which respectively control the importance of $L_{d}$ and $L_{disc}$. The experiments are conducted on an NVIDIA GeForce RTX 3080Ti GPU.
4.2. Performance Evaluations
Here, we compare our method with 5 SOTA detection methods, including LID (lid, ), MD (md, ), Steg (steg, ), SID (SID, ) and SA (txt_advdetection, ).
4.2.1. Generalization performances against unseen adversarial attacks
@@
Tables 3, 4, and 5 present the cross-attack detection results across 7 unseen testing adversarial attacks on CIFAR-10, SVHN, and ImageNet, respectively. Existing adversarial detection methods, which named without ‘-PADs’, are trained on a single attack with their original settings. Each of their results is the averaged detection accuracy with 10 training attacks being utilized one after another in the training process. On the contrary, those with ‘-PADs’ are trained on multiple attacks to be fairly compared to our AED-PADA.
These results provide compelling evidence that our approach obviously outperforms these state-of-the-art adversarial detection methods on the generalization ability. Note that the performance of our approach (99.544%) surpasses that of SID (48.719%) by a significant margin, i.e., 50.825%, under the settings of ImageNet and VGG-16. Besides, our superiority is particularly achieved in challenging scenarios, where the maximum magnitude of the adversarial perturbation is merely set to 2. However, in previous studies, the maximum magnitude is typically set to 4 or 8. The reduction in the magnitude of perturbations significantly increases the difficulty of detecting adversarial examples and reduces the detection performance. The experimental results in Tables 3, 4, and 5 demonstrate that our AED-PADA is effective in the scenarios characterized by subtle adversarial perturbations.
As can be observed, our AED-PADA demonstrates better generalization ability on ImageNet compared to that on CIFAR-10 and SVHN, because the image resolution in ImageNet ($224\times 224$) is higher than that in CIFAR-10 and SVHN ($32\times 32$). Intuitively, a larger image tends to provide a larger space for adversarial perturbation generation, which makes the differences between the perturbations obtained from various attack methods more pronounced, i.e., reducing the overlaps between different adversarial attacks in the adversarial feature space. By selecting PADs, our AED-PADA is able to occupy a larger feature space, thus achieving a better generalization ability.
Besides, PADs may not be suitable to be directly applied to the existing methods. As can be observed, if the SOTA methods directly adopt PADs for training (denoted as X-PADs), their performances may not always increase. For instance, the generalization performance of LID-PADs on CIFAR-10 is inferior to that of LID. We postulate that this discrepancy can be attributed partially to the reduction in the volume of the training data, and partially to the incompatibility between PADs and the framework of LID. This observation further verifies the effectiveness of our PADA in our AED-PADA framework.
4.2.2. Generalization performances across different backbones and datasets
@@
To thoroughly evaluate the generalization ability of adversarial detection methods, it is also vital to assess their performances in the scenarios where the training and testing adversarial examples are from different backbones and datasets. A detection method with better cross-attack, cross-backbone, and cross-dataset capabilities tends to be better suited for practical applications in complex environments.
Table 6 displays the generalization performances of different detection methods across different backbones and datasets. The cross-backbone results are obtained on CIFAR-10, and the cross-dataset results are obtained when the backbone is ResNet-18. In general, results from both scenarios demonstrate that our method exhibits the best generalization ability. Since the structures tend to vary significantly among different types of backbones, which will result in different feature dimensions between training and testing, both LID and MD are not applicable in cross-backbone experiment, i.e., they lack the capability to generalize across backbones, because they utilize the intermediate features of the backbones for detection.
Table 6. Comparison of generalization performances across different backbones and datasets. The bolded and the underlined values represent the best and the second best results for each column, respectively.
@@
Cross-Backbone Cross-Dataset
@@
Table 7. Comparison of generalization performances across different maximum perturbation magnitudes on CIFAR-10 when the backbone is ResNet-18. The maximum magnitude of the testing perturbation is set to 1, 2, 4, 8, respectively.
Detector Known Unseen
2 1 4 8 avg
LID [13] 84.4 56.2 71.0 76.1 67.8
MD [14] 69.6 57.1 66.2 66.4 63.2
Steg [15] 85.2 67.8 90.0 90.4 82.7
SID [16] 76.0 55.2 68.0 64.6 62.6
SA [17] 87.9 75.7 89.8 88.9 84.8
Ours 94.7 89.9 97.0 95.2 94.0
Table 8. Comparison of generalization performances across different maximum perturbation magnitudes on SVHN when the backbone is ResNet-18. The maximum magnitude of the testing perturbation is set to 1, 2, 4, 8, respectively.
Detector Known Unseen
2 1 4 8 avg
LID [13] 69.4 55.0 69.9 75.0 66.6
MD [14] 63.1 55.4 63.4 65.2 61.3
Steg [15] 81.0 58.3 94.7 95.4 82.8
SID [16] 70.5 51.9 64.8 65.4 60.7
SA [17] 74.8 55.7 83.9 84.1 74.6
Ours 93.3 89.1 98.6 99.5 95.7
4.2.3. Generalization performances across different maximum perturbation magnitudes
Here, we evaluate the robustness of the detection models by focusing on a challenging scenario where the maximum perturbation magnitude of the adversarial examples to be detected is unknown. All the detection models are trained on adversarial examples with a maximum perturbation magnitude of 2, while the maximum magnitude of the testing adversarial perturbations is varied across 1, 2, 4 and 8, respectively. Tables 8 and 8 present the generalization performances across different maximum perturbation magnitudes on CIFAR-10 and SVHN, respectively. Based on these results, our AED-PADA consistently outperforms the state-of-the-art adversarial detection methods across all the testing scenarios, demonstrating superior generalization ability of our method against the adversarial perturbations with unseen perturbation magnitudes.
4.3. Effectiveness of AD clustering and PADs Selection
Selecting multiple adversarial attacks from the same cluster leads to redundancy, as these candidate attacks are quite similar. The combination of them covers a small feature space, causing poor generalization. On the other hand, choosing from different clusters effectively avoids this issue. Consequently, as shown in Table 9 and Table 10, we set up two groups of experiments, i.e., selecting ADs from Same Cluster and Cross Cluster, to verify the effectiveness of the AD Clustering and PADs Selection of AED-PADA, against the 7 unseen attacks. ‘Same Cluster’ and ‘Cross Cluster’ respectively represent that the ADs employed as the sources domains are selected from the same AD cluster and different AD clusters.
Table 9 presents the generalization performances on CIFAR-10 and ResNet-18. The result of AD clustering is {FGSM, PGD, DIM, MI-FGSM, SI-NI-FGSM}-{BIM, ILA, YA-ILA}-{C&W, DeepFool}. For ‘Same Cluster’, the best, worst and mean values of averaged results are 90.444%, 93.494%, and 92.150%, respectively. For ‘Cross Cluster’, the best, worst, and mean values are 94.694%, 92.429%, and 93.892%, respectively. Similarly, Table 10 shows the results on CIFAR-10 and VGG-16. The result of AD clustering is {BIM, ILA, YA-ILA, PGD}-{DIM, MI-FGSM, FGSM}-{CW, DeepFool, SI-NI-FGSM}. For ‘Same Cluster’, the best, worst and mean values of the averaged results are 87.132%, 88.019%, and 87.461%, respectively. For ‘Cross Cluster’, the best, worst, and mean values of the averaged values are 88.264%, 87.406%, and 87.818%, respectively.
Table 9. Detection Results of different source domains on CIFAR-10 when the backbone is ResNet-18.
@@ 4.3. Effectiveness of AD clustering and PADs Selection
Same Cluster FGSM+SI-NI-FGSM+PGD 90.444 -
@@
Table 10. Detection Results of different source domains on CIFAR-10 when the backbone is VGG-16.
Origin Detector Averaged Accuracy (%) CEFS($\uparrow$)
Same Cluster BIM+ILA+YA-ILA 87.389 -
BIM+ILA+PGD 87.153 -
BIM+YA-ILA+PGD 87.132 -
ILA+YA-ILA+PGD 87.439 -
DIM+MIM+FGSM 87.634 -
CW+DeepFool+SI-NI-FGSM 88.019 -
Cross Cluster ILA+DIM+CW 88.264 18.692
PGD+DIM+CW 88.236 18.605
ILA+FGSM+CW 87.969 18.532
YA-ILA+FGSM+CW 87.897 18.527
YA-ILA+MI-FGSM+SI-NI-FGSM 87.670 18.518
BIM+DIM+SI-NI-FGSM 87.676 18.410
YA-ILA+FGSM+SI-NI-FGSM 87.425 18.388
YA-ILA+DIM+SI-NI-FGSM 87.406 18.385
Based on the results, we can obtain three observations. Firstly, the mean results of ‘Cross Cluster’ selection is higher than of ‘Same Cluster’ selection, and the best result of ‘Cross Cluster’ selection is also superior. Specially, On the setting of CIFAR-10 and ResNet-18, the mean results of ‘Cross Cluster’ selection is even higher than of the best results of ‘Same Cluster’ selection. This verifies the effectiveness of our AD Clustering. Apparently, source domains from ‘Cross Cluster’ can certainly enhance the generalization ability of the detection methods.
Secondly, although ‘Cross Cluster’ selection in general achieves better results, it cannot guarantee that the randomly selected ADs (from ‘Cross Cluster’s) always give better performance than the results of ‘Same Cluster’ selected ADs. When the PADs with a very low CEFS score are employed as the source domains for detection, the generalization performance may be worse than that of ‘Same Cluster’ selection. This observation further verifies the effectiveness of our PADs Selection.
Thirdly, as shown in the ‘Cross Cluster’ part of table 9 and table 10, for the majority of results, a higher CEFS value of PADs induces a higher averaged accuracy of the proposed detection. This observation suggests that the proposed CEFS is a proper guide for selecting PADs, i.e., a higher CEFS score indicates that the corresponding PADs can cover a larger proportion of the entire feature space, and the PADA stage, which uses the PADs as source domains, gives a better generalization performance.
4.4. The effect of different numbers of clusters
[…]
As can be observed, our automatic CH score based method can yield superior performance on SVHN with VGG-16 being the backbone, and give comparable performance to the best result on other settings. Although the best result with manually selecting $K$ exceeds our automatic method for 0.2%-0.3% in terms of the averaged accuracy, it requires 2.3×2.3$\times$2.3 × more parameters and gives a 3×3$\times$3 × slower speed. Moreover, in real-world scenarios, the adversarial attacks in the training set may far exceed 10 types, and it is clearly unwise to test the detection performance of each clustering result individually. Consequently, this suggests that our method actually achieves a better balance between the training costs and generalization performances.
Table 11. Detection Performance across seven unseen testing adversarial attacks with different number of clusters $K$.
@@ 4.4. The effect of different numbers of clusters
CIFAR-10 (ResNet-18) 91.936 $94.694^{{\dagger}}$ 94.784 94.571 94.561 95.018
CIFAR-10 (VGG-16) 87.946 $88.264^{{\dagger}}$ 87.964 87.991 88.104 88.498
SVHN (ResNet-18) 91.497 $94.344^{{\dagger}}$ 91.257 91.726 94.999 94.801
SVHN (VGG-16) 89.183 89.269 $\textbf{91.219}^{{\dagger}}$ 90.056 90.026 90.713
Table 12. Detection Performance of different feature extractions in PADA of our AED-PADA.
@@
4.5. Effectiveness of the adversarial feature enhancement in PADA